OpenStudy (anonymous):

which is not an output for 2x squared-5? -3 45 27 -8

OpenStudy (anonymous):

the in/out table is obtained by the equation y = 2x^2 - 5. so if you add 5 and divide by 2 you'll get. (y+5)/2 = x^2. now, each of those are the y's (the output of the original equation) so if you stick in those values for y and you don't get a perfect square, then that choice is correct answer. for example, y=-3. x^2=1 perfect square. do this with the other numbers... until you don't get a perfect square.

OpenStudy (anonymous):

He got the best medal^

OpenStudy (anonymous):

you have the answer?

OpenStudy (anonymous):

-8

OpenStudy (anonymous):

I was typing it our but you had the best answer :)

OpenStudy (anonymous):

good!... i thought what i wrote didn't make any sense... :)

OpenStudy (anonymous):

out*

OpenStudy (anonymous):

thanks Abbie23... :)

OpenStudy (anonymous):

You're welcome dpalnc! :)

OpenStudy (anonymous):

@Abbie23 , you could also put your solution down so you get a medal from me.

OpenStudy (anonymous):

just did.. besides, there's more than 1 way to look at a problem. if the asker could not understand my method, maybe she'll understand yours... :)